Reece:

Traveling with the current, a certain boat went 15 km/h. Against the same current, it only went 1 km/h. Find the speed of the boat in still water and the speed of the current. The speed of the boat is? The speed of the current is?

Reece:

i need help guys plz

Vocaloid:

Let the speed of the boat be b Let the speed of the current be c Going with the current: speed = b + c = 15 Against the current: b - c = 1 Solve for b

Reece:

but whats the answer

Vocaloid:

Please be mindful that we don’t give direct answers here To solve the system, start by adding the two equations together

Reece:

ok ;(